# SReferences

### v2

@@ II Erasure and error rates of Logical fusions
For REP logical fusions, as outlined in the main text, $m$ fusions are attempted ballistically between pairs of photons in the logical qubits. All $m$ physical $XX$ outcomes are required to obtain the $\overline{XX}$ outcome, whereas only a single physical $ZZ$ outcome is required for $\overline{ZZ}$. The corresponding recovery probabilities are then
@@
The logical error rates of REP fusion can be found by counting the number of physical fusion errors and determining whether this gives a logical error. Let us assume that $X$ and $Z$ errors occur on each fusion outcome with probabilities $p_{X}$ and $p_{Z}$, which flip $ZZ$ and $XX$ outcomes, respectively. If the probability that a photon acquiring $Z$ error prior to fusion is $e_{Z}$, then the probability of having a flipped $XX$ fusion outcome is the probability of odd numbers of photon $Z$ error which is $p_{X}=2e_{Z}$. Similarly on the logical level, since the $Z$ repetition code offers no assistance to the recovery of $\overline{XX}$ outcome, the probability of a logical error in $\overline{XX}$ is the probability of having an odd number of $Z$ fusion errors in $m$ physical fusion attempts:
@@
With majority voting, logical flips in the $\overline{ZZ}$ outcome occur when half or more of the $t$ recovered physical $ZZ$ outcomes have an error.
@@
These expressions denote the conditional probability of an error given that the logical outcomes were recovered. We can improve the error tolerance by treating a $\overline{ZZ}$ outcome as erased when there are equal numbers of good and erroneous physical $ZZ$ outcomes, as we have no confidence in the value of the $\overline{ZZ}$ operator.
RUS logical fusions have the same requirements for obtaining logical fusion outcomes, but now the code size can vary. Moreover, only a single $ZZ$ operator is required to recover $\overline{ZZ}$ since RUS terminates further fusion attempts once $ZZ$ is obtained. Four distinct events can occur after a logical RUS fusion: ① $0\leq j<N$ failures occur before a successful fusion, recovering both $\overline{XX}$ and $\overline{ZZ}$. ②: $N$ consecutive failures occur, recovering only $\overline{XX}$. ③: $20\leq j\leq N-$ 2 consecutive failures occur, followed by a loss in the $(j+1)$-th fusion. The physical fusion failure basis is changed to $ZZ$. There are then $20\leq k<N-j-$ 2 additional losses before a single non-lossy fusion that recovers $\overline{ZZ}$ only. ④: As for ③, but all fusions after the $j$ failures have loss. No outcomes are recovered, corresponding to logical erasure. These events occur with probabilities
@@
where $P_{1}+P_{2}+P_{3}+P_{4}=1$. The recovery probabilities for the respective outcomes are then $R_{\overline{XX}}=P_{1}+P_{2}$, and $R_{\overline{ZZ}}=P_{1}+P_{3}$. In the simulation, we draw samples of RUS logical fusion outcomes from the weighted probability distribution $\{P_{1},P_{2},P_{3},P_{4}\}$ using Eq. (S12).
The logical fusion error rates are calculated similarly to before. However, since $\overline{ZZ}$ is always recovered by one $ZZ$ outcome, this means that majority voting is no longer possible, thus the logical fusion error rate is equal to the physical fusion error rate.
@@
Figure S4: Monte-Carlo simulation of RUS fusion. (a) Simulated logical fusion success probability as a function of end-to-end efficiency $\eta$. The simulation uses $N=11$ with $10^{4}$ trials. Analytical curves are obtained using Eq. (S12). (b) Distribution of number of fusion attempts $m$ in a RUS logical fusion. $z$-axis is the probability that the RUS fusion stops at $m$-th attempt. At $20\%$ loss, there is a $84\%$ chance that the RUS fusion terminates within $m=3$ attempts.
@@ III RUS fusion with re-initialization of emitters
Figure S5: Schematics of RUS fusion with re-attempts. A RUS fusion (blue) at time $t=1$ can be re-attempted when the two emitters are previously disconnected from the graph due to two logical fusions (pink) resulting in biased $\overline{ZZ}$ outcome at $t=0$.
@@
For RUS logical fusions, when a photon is lost during a physical fusion, the $XX$ outcome is irreversibly lost thus $\overline{XX}$ cannot be recovered. On the other hand, the $\overline{ZZ}$ outcome can still be obtained by retrieving $ZZ$ in the next fusion attempt (Sec. VI.2), due to re-emission of the photons followed by single-qubit measurement in the Pauli-$Z$ basis. Crucially, when the $ZZ$ outcome is retrieved, the two emitter spins are indirectly measured in the Pauli-$Z$ basis, since the measured photon and emitter share the stabilizer $ZZ$. This implies that once the logical fusion results in $\overline{ZZ}$ outcome only, the emitter spins become disconnected from the original graph with their states effectively re-initialized.
@@
In the case of two neighboring logical fusions being biased in $\overline{ZZ}$ at an initial time $t=0$ (Fig. S5), both emitters at $t=1$ are re-initialized and detached from any previous correlations, allowing the next logical fusion to attempt RUS as many times as possible. The probability of recovering both $\overline{XX}$ and $\overline{ZZ}$ is therefore drastically increased. For realistic simulation, this number is set to be equal to the number of maximum physical fusion attempts $N$. In the absence of photon loss, the probability of recovering both logical fusion outcomes scales as $1-2^{-N}$ thus making the logical fusion near deterministic.
@@
Interestingly, in a realistic experiment this trick can be applied to the first layer of logical fusions in the FFCC lattice to make them near deterministic, since all emitters are initially disconnected. Here for simulations we are only interested in the bulk of the lattice thus it is applied only to the subsequent layers. It is important to note that this scheme relies on the previous two neighboring logical fusions only retrieving the $\overline{ZZ}$, which is a second-order effect, and thus, this additional trick gives a modest improvement of the loss threshold to 8%.

### v3

@@ II Erasure and error rates of encoded fusions
For REP encoded fusions, as outlined in the main text, $m$ fusions are attempted ballistically between pairs of photons in the encoded qubits. All $m$ physical $XX$ outcomes are required to obtain the $\overline{XX}$ outcome, whereas only a single physical $ZZ$ outcome is required for $\overline{ZZ}$. The corresponding recovery probabilities are then
@@
The logical error rates of REP fusion can be found by counting the number of physical fusion errors and determining whether this gives a logical error. Let us assume that $X$ and $Z$ errors occur on each fusion outcome with probabilities $p_{X}$ and $p_{Z}$, which flip $ZZ$ and $XX$ outcomes, respectively. If the probability that a photon acquiring $Z$ error prior to fusion is $e_{Z}$, then the probability of having a flipped $XX$ fusion outcome is the probability of odd numbers of photon $Z$ error which is $p_{Z}=2e_{Z}$. Similarly on the logical level, since the $Z$ repetition code offers no assistance to the recovery of $\overline{XX}$ outcome, the probability of a logical error in $\overline{XX}$ is the probability of having an odd number of $Z$ fusion errors in $m$ physical fusion attempts:
@@
With majority voting, flips in the $\overline{ZZ}$ outcome occur when half or more of the $t$ recovered physical $ZZ$ outcomes have an error.
@@
These expressions denote the conditional probability of an error given that $\overline{XX}$ and $\overline{ZZ}$ were recovered. We can improve the error tolerance by treating a $\overline{ZZ}$ outcome as erased when there are equal numbers of good and erroneous physical $ZZ$ outcomes, as we have no confidence in the value of the $\overline{ZZ}$ operator.
RUS encoded fusions have the same requirements for obtaining encoded fusion outcomes, but now the code size can vary. Moreover, only a single $ZZ$ operator is required to recover $\overline{ZZ}$ since RUS terminates further fusion attempts once $ZZ$ is obtained. Four distinct events can occur after a logical RUS fusion: ① $0\leq j<N$ failures occur before a successful fusion, recovering both $\overline{XX}$ and $\overline{ZZ}$. ②: $N$ consecutive failures occur, recovering only $\overline{XX}$. ③: $20\leq j\leq N-$ 2 consecutive failures occur, followed by a loss in the $(j+1)$-th fusion. The physical fusion failure basis is changed to $ZZ$. There are then $20\leq k<N-j-$ 2 additional losses before a single non-lossy fusion that recovers $\overline{ZZ}$ only. ④: As for ③, but all fusions after the $j$ failures have loss. No outcomes are recovered, corresponding to logical erasure. These events occur with probabilities
@@
where $P_{1}+P_{2}+P_{3}+P_{4}=1$. The recovery probabilities for the respective outcomes are then $R_{\overline{XX}}=P_{1}+P_{2}$, and $R_{\overline{ZZ}}=P_{1}+P_{3}$. In the simulation, we draw samples of RUS encoded fusion outcomes from the weighted probability distribution $\{P_{1},P_{2},P_{3},P_{4}\}$ using Eq. (S12).
The encoded fusion error rates are calculated similarly to before. However, since $\overline{ZZ}$ is always recovered by one $ZZ$ outcome, this means that majority voting is no longer possible, thus the encoded fusion error rate is equal to the physical fusion error rate.
@@
Figure S4: Monte-Carlo simulation of RUS fusion. (a) Simulated encoded fusion success probability as a function of end-to-end efficiency $\eta$. The simulation uses $N=11$ with $10^{4}$ trials. Analytical curves are obtained using Eq. (S12). (b) Distribution of number of fusion attempts $m$ in a RUS encoded fusion. $z$-axis is the probability that the RUS fusion stops at $m$-th attempt. At $20\%$ loss, there is a $84\%$ chance that the RUS fusion terminates within $m=3$ attempts.
@@ III RUS fusion with re-initialization of emitters
Figure S5: Schematics of RUS fusion with re-attempts. A RUS fusion (blue) at time $t=1$ can be re-attempted when the two emitters are previously disconnected from the graph due to two encoded fusions (pink) resulting in biased $\overline{ZZ}$ outcome at $t=0$.
@@
For RUS encoded fusions, when a photon is lost during a physical fusion, the $XX$ outcome is irreversibly lost thus $\overline{XX}$ cannot be recovered. On the other hand, the $\overline{ZZ}$ outcome can still be obtained by retrieving $ZZ$ in the next fusion attempt (Sec. VI.2), due to re-emission of the photons followed by single-qubit measurement in the Pauli-$Z$ basis. Crucially, when $ZZ$ outcomes are retrieved from two neighboring encoded fusions, the two emitter spins involved are indirectly measured in the Pauli-$Z$ basis, since the measured photon and its entangled emitter share the stabilizer $ZZ$. This implies that once the encoded fusion results in $\overline{ZZ}$ outcome only, the emitter spins become disconnected from the original graph with their states effectively re-initialized.
@@
In the case of two neighboring encoded fusions being biased in $\overline{ZZ}$ at an initial time $t=0$ (Fig. S5), both emitters at $t=1$ are re-initialized and detached from any previous correlations, allowing the next encoded fusion to attempt RUS as many times as possible. The probability of recovering both $\overline{XX}$ and $\overline{ZZ}$ is therefore drastically increased. For realistic simulation, this number is set to be equal to the number of maximum physical fusion attempts $N$. In the absence of photon loss, the probability of recovering both encoded fusion outcomes scales as $1-2^{-N}$ thus making the encoded fusion near deterministic.
@@
Interestingly, in a realistic experiment this trick can be applied to the first layer of encoded fusions in the FFCC lattice to make them near deterministic, since all emitters are initially disconnected. Here for simulations we are only interested in the bulk of the lattice thus it is applied only to the subsequent layers. It is important to note that this scheme relies on the previous two neighboring encoded fusions only retrieving the $\overline{ZZ}$, which is a second-order effect, and thus, this additional trick gives a modest improvement of the loss threshold to 8%.
IV Fault-tolerant region with multiple errors
A practical quantum computer will likely be simultaneously subject to multiple error channels, it is therefore useful to know whether a given set of errors lies within the fault-tolerant region of the quantum computing architecture.
In Fig. S6, we simulate the combined effect of photon loss, distinguishability between emitters, and spin depolarizing errors, and present the region (shaded in pink) in which the logical error can be corrected. The colored curves at the boundaries are fault-tolerant thresholds obtained with either two of the above errors. Here we consider the synchronous FFCC architecture, using RUS encoded fusions in which the re-initialization of emitters is allowed for maximum loss threshold. Both the numbers of maximum physical fusion attempts and RUS re-attempts are set to $N=10$ since the thresholds shown in the main text begin to saturate at $N\geq 10$.
Figure S6: Fault-tolerant region of the sFFCC architecture, using RUS fusions with emitter re-initialization. We consider photon loss, distinguishability between emitters, and spin depolarizing error. The number of maximum fusion attempts and the number of RUS re-attempts are both set to $N=10$.
The noise simulation is a two-step process, similar to the one with only spin noise discussed in Sec. VII, which begins by first sampling encoded fusion outcomes under the presence of both photon loss and distinguishability from the weighted probability distribution $\{P^{\bar{\text{X}}\bar{\text{Z}}}_{N},P^{\bar{\text{X}}}_{N},P^{\bar{\text{Z}}}_{N},P^{\bar{\text{e}}}_{N}\}$, where the analytical formula for each probability is:
$\overline{XX}\text{ and }\overline{ZZ}:\quad\quad P^{\bar{\text{X}}\bar{\text{Z}}}_{N}$ $=P_{\text{s}}+P_{\text{s}}\sum^{N-1}_{i=1}(P_{\text{f}})^{i};$
$\overline{XX}\text{ only }:\quad\quad P^{\bar{\text{X}}}_{N}$ $=(P_{\text{f}})^{N}=(\eta^{2}P^{\text{e}}_{Z}P_{X})^{N};$
$\text{Erasure }:\quad\quad P^{\bar{\text{e}}}_{N}$ $=(P_{l}+P^{\text{e}}_{\text{f}})(P_{l})^{N-1}+(P_{l}+P^{\text{e}}_{\text{f}})\sum^{N-1}_{i=1}(P_{\text{f}})^{i}(P_{l})^{N-1-i};$
$\overline{ZZ}\text{ only }:\quad\quad P^{\bar{\text{Z}}}_{N}$ $=1-P^{\bar{\text{X}}}_{N}-P^{\bar{\text{e}}}_{N}-P^{\bar{\text{X}}\bar{\text{Z}}}_{N}.$ (S14)
$P^{j}_{N}$ is the probability that an encoded fusion results in measurement outcome $j$ within $N$ physical fusion attempts. With finite photon distinguishability $1-V\neq 0$, a physical fusion could in principle result in erasure of both outcomes, i.e., when a $\{ZZ,XX\}$ fusion fails without recovering $XX$. In such a case, the $ZZ$/$XX$ outcome is erased with probabilities $P^{\text{e}}_{Z}=\frac{1}{2}=1-P_{Z}$ and $P^{\text{e}}_{X}=\frac{1-V}{4}=1-P_{X}$. Including photon loss, there are in total five possible fusion outcomes: (1) successful fusion without erasure in $XX$, denoted by $P_{\text{s}}$, (2) successful fusion with $XX$ erasure $P^{\text{e}}_{\text{s}}$, (3) failed fusion without $XX$ erasure $P_{\text{f}}$, (4) failed fusion with $XX$ erasure $P^{\text{e}}_{\text{f}}$, and (5) fusion erasure $P_{l}$ owing to photon loss. For the first four cases, both photons are not lost. As a result, the probability for each of the above five cases is:
$P_{\text{s}}$ $=\eta^{2}P_{Z}P_{X};\quad\quad P^{\text{e}}_{\text{s}}=\eta^{2}P_{Z}P^{\text{e}}_{X};\quad\quad P_{\text{f}}=\eta^{2}P^{\text{e}}_{Z}P_{X};\quad\quad P^{\text{e}}_{\text{f}}=\eta^{2}P^{\text{e}}_{Z}P^{\text{e}}_{X};\quad\quad P_{l}=1-\eta^{2}.$ (S15)
After sampling a list of outcomes for all encoded fusions, a search is performed over each layer of the lattice to locate encoded fusions that end up only in $\overline{ZZ}$. If the condition for RUS emitter re-initialization is fulfilled (Supplementary Note III): two neighboring encoded fusions are biased in $\overline{ZZ}$ at layer $i$, and the encoded fusion at layer $i+1$ only gives either one or no measurement outcome, then its outcome is re-sampled for $N$ additional RUS re-attempts.
Once a final list of encoded fusion outcomes is obtained, a fusion error sampling process for distinguishability and spin depolarizing noise begins for each encoded fusion:
  1. 1.
    If RUS fusion results in both $\overline{XX}$ and $\overline{ZZ}$, the $\overline{ZZ}=ZZ$ outcome has a probability of $\frac{1-V}{2}$ being flipped due to photon distinguishability. This $ZZ$ outcome would have resulted from the last physical fusion attempt $m\leq N$, since encoded fusion stops once $\overline{ZZ}$ is recovered. $m$ is sampled from a list of $N$ probabilities $\mathcal{P}_{\bar{X}\bar{Z}}=\{P_{i}\}_{i\in[1,N]}$, where $P_{i}=P^{\bar{\text{X}}\bar{\text{Z}}}_{i}-P^{\bar{\text{X}}\bar{\text{Z}}}_{i-1}$ is the probability that RUS terminates on exactly the $i$-th attempt. Note $P^{\bar{\text{X}}\bar{\text{Z}}}_{0}=0$. The value of $m$ allows one to extract the error on the $m$-th photon of an encoded qubit in the resource state. Odd numbers of $X$ ($Z$) photon error(s) would lead to a flip in $\overline{ZZ}$ ($\overline{XX}$).
  2. 2.
    If RUS fusion ends up with only $\overline{ZZ}$, $m$ is sampled from the list $\mathcal{P}_{\bar{Z}}=\{P_{i}\}_{i\in[1,N]}$ where $P_{i}=P^{\bar{\text{Z}}}_{i}-P^{\bar{\text{Z}}}_{i-1}$. Note $P^{\bar{\text{Z}}}_{0}=0$. Errors in the photons involved in the $m$-th physical fusion would contribute to an error in $\overline{ZZ}$.
  3. 3.
    When a RUS fusion leads to $\overline{XX}$ only, $m$ is sampled from $\mathcal{P}_{\bar{X}}=\{P^{\bar{X}}_{i}\}_{i\in[1,N]}$. In such case, $\overline{XX}=(XX)^{\otimes m}$, and hence an odd number of errors in the first $m$ photons in each of the two encoded qubits would flip $\overline{XX}$.
V Comparing theoretical thresholds with current state-of-the-arts
The threshold analysis presented in the main text is hardware-agnostic, as it focuses on error channels intrinsic to quantum emitters. Generally, any quantum emitter with access to spin and optical transitions (key prerequisites for the deterministic generation of entangled-photon cluster states) is inherently susceptible to spin dephasing, photon loss, and distinguishability errors. The fault-tolerant thresholds we have derived aim to provide a useful baseline, which helps consolidate experimental efforts towards improving specific areas crucial for practical implementation of the architecture. In this context, it is instructive to compare the current state-of-the-art figures of merits with the theoretical thresholds, as it highlights areas which require further advancements.
Architecture End-to-end efficiency Emitter-emitter indistinguishability Required $T_{2}$ (photon generation time $\tau$)
sFFCC, RUS ($N=10$) $92\%$ $96\%$ $500\tau$
Emitter platform Outcoupling Emitter-emitter $T_{2}$ time
efficiency indistinguishability Required Achieved
Neutral atom in optical cavity $66\%$ [5] $95.5\%$ [45] 37.537.537.5~{}37.5ms ($\tau=75~{}\mu$[37]) $1.2$ ms [5]
Quantum dot $57\%$ [46], $71\%$ [35] $93\%$ [32] $34.5~{}\mu$s ($\tau=68.9$ ns [8]) $113~{}\mu$[43]
SiV in diamond cavity $42\%$ [47] $72\%$ [48] $0.071$ ms ($\tau=142$ ns [47]) $>1.5~{}$ms [49], $13$ ms [50]
Table S1: Comparison between experimental state-of-the-arts and thresholds required for sFFCC. Upper panel: required specifications to reach fault-tolerance for sFFCC. It is important to note that these parameters are derived from the relevant thresholds assuming other errors are absent. Lower panel: Current state-of-the-arts. Outcoupling efficiency is considered to allow comparison between different material platforms. Emitter-emitter indistinguishability is defined by the measured raw Hong-Ou-Mandel visibility between photons from two emitters.
The upper panel of Table S1 shows the required figures of merits derived from thresholds for the sFFCC architecture with RUS fusions and emitter re-initialization. The lower panel lists the experimental state-of-the-art parameters for three prominent material platforms, which have emitters with a built-in spin echo sequence. Note that the table is not meant to be exhaustive, and some numbers may not be directly comparable due to different experimental conditions. From the table, several conclusions could be drawn: (1) photon loss is the primary bottleneck for FBQC. The $92\%$ required end-to-end efficiency derived from photon loss threshold includes: the outcoupling efficiency of each resource state generator, fiber transmission, coupling efficiency to chip, and detection efficiency etc. While tremendous progress has been made in improving the efficiencies of integrated components [36], the outcoupling efficiency still requires significant advancements to bridge the gap. (2) Both single neutral atom and quantum dot systems have demonstrated close-to-threshold photon distinguishability between emitters. (3) Both quantum dot and SiV platforms may have reached the required tolerance for spin decoherence, while single neutral atom suffers from long generation time $\tau=75~{}\mu$s between two consecutively emitted photons owing to both qubit transfer and rotation pulses [37]. However, for both quantum dot and SiV, additional decoherence processes like laser-induced or heating-induced dephasing, should be considered in future simulations to obtain a more realistic threshold.